OpenStudy (anonymous):

for f(x)= 1/x^2 and g(x)=x^2. find (fog)(x)

OpenStudy (tkhunny):

Isn't that just f(g(x))? Let's see what you get. Do what the notation suggests.

OpenStudy (anonymous):

i just plug it in?

OpenStudy (tkhunny):

It's called "substitution", but that's the idea.

OpenStudy (anonymous):

oh okay well , would i plug or substitute the x^2 into the other x^2?

OpenStudy (anonymous):

\[(fog)(x)=\frac{1}{x^4}\]

OpenStudy (tkhunny):

@electrokid Nover do that again. If h(x) = 3x + 2 h(2) = 3(2) + 2 h(\(\sqrt{2}) = 3\sqrt{2} + 2\) h(frog) = 3(frog) + 2 h(g(x)) = 3(g(x)) = 2 Whatever it is, just substiute it.

OpenStudy (anonymous):

okay but for this equation how would i do that if they are both x^2?

OpenStudy (tkhunny):

The are NOT both \(x^{2}\). They are as they are defined. \(f(x) = \dfrac{1}{x^{2}}\;and\;g(x) = x^{2}\) \(f(g(x)) = \dfrac{1}{[g(x)]^{2}} = \dfrac{1}{(x^{2})^{2}} = \dfrac{1}{x^{4}}\) Whatever it is, just substitute it. No need to overthink it. Just do it!

OpenStudy (anonymous):

ohh okay . thanks